OpenStudy (anonymous):

I JUST CAN'T get the hang of these. In tears! Two cyclists leave towns 76 mi apart at the same time and travel toward each other. One cyclist travels 6 mph faster than the other. If they meet in 2 hours, what is the rate of each cyclist?

OpenStudy (zpupster):

what is rate of speed of cyclist 1?

OpenStudy (anonymous):

umm... I need to figure the rate of both bikes

OpenStudy (zpupster):

well the answer is i do not know so we would let that be x cyclist 2 would x+6 you know the time 2 and the distance 76 2x+2(X+6) = 76 solve this for x

OpenStudy (jhannybean):

You could also think about it like this: since there are 2 cyclists and 1 is traveling FASTER than the other, then let: cyclist 1 = x (the one that's moving slower) cyclist 2 = x+6 (this guy is travelling 6 miles FASTER than the first guy)

OpenStudy (anonymous):

2x + 2(x+6) = 76 2x + 12 = 76

OpenStudy (zpupster):

2x + 2(x+6) = 76 2x +2x+ 12 = 76

OpenStudy (anonymous):

2x + 2(x+6) = 76 2x + 12 = 76 12/76 = 6.333

OpenStudy (jhannybean):

As d = rt (distance = rate TIMES time) , and we know the rate of cyclist 1 and 2, then : d\(_1\) = 70 mi = 2x d\(_2\) = 2(x+6)

OpenStudy (jhannybean):

Combining these two, since they are both travelling the same distance, we get \[70 = 2x +2(x+6)\]

OpenStudy (anonymous):

2x + 2(x+6) = 76 4x + 24 = 76

OpenStudy (jhannybean):

First of all, is what me and @zpupster making sense?

OpenStudy (anonymous):

4x + 12 = 76

OpenStudy (jhannybean):

It's easy to plug and chug, but are you understanding the concept?

OpenStudy (anonymous):

I think it is a rate x distance = time?

OpenStudy (jhannybean):

No. distance = rate x time

OpenStudy (jhannybean):

That is why you are multiplying `2` to both rates,

OpenStudy (jhannybean):

2(x) and 2(x+6)

OpenStudy (anonymous):

6 x 2 = Distance

OpenStudy (jhannybean):

Does 6 \(\times\) 2 = 70?

OpenStudy (anonymous):

No 12

OpenStudy (zpupster):

this was right just solve for x 4x + 12 = 76

OpenStudy (anonymous):

x = 76 - 12

OpenStudy (jhannybean):

No.

OpenStudy (jhannybean):

4x = 76 - 12*

OpenStudy (zpupster):

4x = 76 - 12

OpenStudy (anonymous):

x = 76 + 12 x = 88

OpenStudy (anonymous):

4/88=11

OpenStudy (jhannybean):

That is wrogn.

OpenStudy (anonymous):

x = 11 11+ 6 = 17 faster

OpenStudy (jhannybean):

You don't really seem like you know what you are doing.

OpenStudy (anonymous):

I have been trying for a week. My teacher is ignoring me

OpenStudy (jhannybean):

First of all, where did your 4 disappear to? 4x = 76 - 12 suddenly became x = 76 - 12

OpenStudy (zpupster):

slow down faine take a breath then let start with 4x = 76 - 12

OpenStudy (anonymous):

That is why I am crying. I have been sitting here all day doing one after another and all were wrong

OpenStudy (anonymous):

I divided my 4 into the 88 to get 11

OpenStudy (jhannybean):

I think what you are confused with starting from the top is understanding what the RATE of each cyclist is, and why one is labeled as x and the other is labeled as x+6.

OpenStudy (anonymous):

x = 64 @zpupster

OpenStudy (jhannybean):

Once you understand that, you can associate that with the distance formula, d = rt, where the time for BOTH is 2 hours.

OpenStudy (anonymous):

4x=64

OpenStudy (zpupster):

yes

OpenStudy (anonymous):

4/64 = 16

OpenStudy (jhannybean):

That's correct.

OpenStudy (zpupster):

divide both sides by 4

OpenStudy (jhannybean):

Now you've got the rate of the slower cyclist.

OpenStudy (anonymous):

16 + 6 = 22

OpenStudy (jhannybean):

All you need to do is figure out how much faster the other one is moving, which you just got as well.

OpenStudy (zpupster):

good job!!!

OpenStudy (jhannybean):

Now retry this problem without looking.

OpenStudy (jhannybean):

Answer me this, what if the problem said `one cyclist is moving SLOWER than the other`?

OpenStudy (jhannybean):

w would you write the rate of cyclist #2?

OpenStudy (anonymous):

How do I medal all of you? I can not believe my stupidity is this obvious and I am feeling dumb. I'm in friggin college for sanity sake! How did I get this far. THANK YOU ALL!!!!!!!

OpenStudy (anonymous):

16 - 6 = 10 If it were 6 mph slower?

OpenStudy (jhannybean):

Awesome.

OpenStudy (anonymous):

I hope these stay in my abccount so I can look back on what we did and try a few alone

OpenStudy (jhannybean):

And if the problem said they were moving against eachother, then you would subtract the rates.